# SpiderGen:Towards Procedure Generation For Carbon Life Cycle Assessments with Generative AI

### v1

@@ 6 Evaluation Results
We present an evaluation of SpiderGen, highlighting key results.
@@
o1-preview $0.051\pm 0.005$ US\$ $4.7\pm 0.28$
o1-mini $0.046\pm 0.006$ US\$ $0.36\pm 0.03$
gpt-4o $0.039\pm 0.003$ US\$ $0.64\pm 0.14$
Table 2: Average PMI Score and cost for SpiderGen using different LLMs. SpiderGen performs better with models, such as OpenAI’s o1-preview and o1-mini. o1-mini provides a good tradeoff with good performance and lowest cost.
@@
We implemented SpiderGen using 3 OpenAI models: gpt-4o, o1-mini, and o1-preview. o1-mini and o1-preview are reasoning models, with o1-mini being a smaller, faster, and cheaper model than the other two. Table 2 compares the performance of these models in terms of the average normalized PMI and cost for generating a single PFG. The two reasoning models (o1-mini and o1-preview) outperform gpt-4o since they are able to provide more detailed outputs. However, o1-mini is close in terms of performance to o1-preview while being 1/13th the cost at less than US \$1 on average per PFG. This cost is far lower than the typical production of a PFG, which can be upwards of US \$25000.
@@
We evaluate the effect of producing a different number of sample products in the first step of SpiderGen (shown in Figure 3 for four product categories). We evaluate the production of PFGs using 3, 6, 9, 12, 15 sample products for each product category. As shown in the figure, the PMI often increases with the addition of more sample products. We note that less “niche” product categories, such as “Grain Mills” benefit from a larger number of sample products, and more “niche” categories, such as “Moka Coffee” benefit from smaller numbers of sample products, and can even have deteriorated PFGs, as the number of sample products increases. Since generating 15 products works well for many scenarios, we use it for our implementation of SpiderGen.
@@
To evaluate SpiderGen end-to-end, we execute our pipeline for 66 product categories and compare the generated PFGs to the ground truth graphs extracted from real LCA documents (EPD).
@@
We compared SpiderGen with the two baseline methods LLMDirect and LLMExample across 40 product categories. Our resuts show that SpiderGen exceeds both baseline methods, providing a median normalized PMI score of $0.043\pm 0.025$ across different product categories as compared to $0.026\pm 0.022$ for LLMDirect and $0.024\pm 0.024$ PMI for LLMExample. Based on our qualitative analysis of a subset of ten products, we find that SpiderGen avoids generating overly specific processes and captures more non-obvious auxiliary processes as compared to the baselines.
Limitations of LLMDirect: Although a Chain-of-Thought process may work for many problems, we found that an algorithm such as LLMDirect is insufficient to solve our problem. Firstly, it is unable to capture all the nuanced steps that are necessary for complete PFGs. LLMDirect results frequently misses steps that are peripheral to the main manufacturing processes. Further, it is unclear whether the resultant PFG is truly applicable to all products in a specific product category and is likely to be overly specific to one subcategory of the product category.
Limitations of LLMExample: While LLMExample is more effective than LLMDirect, since it is able to capture the nuanced steps of the PFG by following an example, it is still insufficient to solve our problem. LLMExample misses peripheral steps that may be important to a broader product category. Additionally, the PFG is still overly specific to a subset of items of the product category. This is particularly likely if the example product category is more niche than the product category information that we are generating.
@@ 7 Discussion and Conclusion
In this work, we introduce SpiderGen, a novel machine learning workflow that automates the generation of PFGs, and introduce evaluation methods for measuring the correctness of PFGs. We present an implementation of SpiderGen, and provide an evaluation of SpiderGen across a set of 66 data points. We find that SpiderGen is able to capture a large portion of required processes for tracking the carbon footprint of various product categories.
@@
Increasing transparency in LCA automation. Recent work indicates that providing transparency in automated LCA methods is a key criteria (ulissi_criteria_2025). Thus, in the design of SpiderGen, we enable transparency by providing sample products, evidence and rationales for processes, and deterministic clustering techniques for forming coarser processes. We hope that this enables human LCA experts to gain a better understanding of how SpiderGen generates PFGs. However, more work must be done to achieve other evaluation goals, such as measuring and indicating the uncertainty of the PFG generation, as well as enabling LCA experts to collaboratively improve the PFG estimates. We hope to explore these possibilities in future work.
Enhancing System Boundaries of PFG Generation. We note that a primary limitation of SpiderGen is derived from the scope, or “system boundary”, of the PFG being in conflict with the system boundary defined appropriately for the product category. In the real world, the system boundary is human-defined. However, for our experiments, we made assumptions about the system boundaries based on ISO standards and encoded them into SpiderGen. We note that this is the root cause for SpiderGen missing processes or hallucinating processes. Although the precision of SpiderGen is relatively high (on average 83%), we note that SpiderGen occasionally produces erroneous processes due to sticking closely to system boundaries that may be inappropriate for a given product. For example, a more common error was to include processes such as “replacement”, “quality control”, or “energy consumption in use” when it was not relevant for a specific product category. We observe in our experiments that SpiderGen misses some the auxiliary processes which are not directly related to the product manufacturing, material composition or usage. We believe that exploring human-AI collaboration methods to make appropriate choices for system boundaries is a promising next step for future work.
@@ Acknowledgments
This work is supported by the National Science Foundation Award CNS-2106299, as well as the National Science Foundation Graduate Research Fellowship Program under Grant No(s) (DGE-2140739). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation).

### v2

@@ 6 Evaluation Results
We present an evaluation of SpiderGen, highlighting key results. Additional evaluation results can be found in the extended version of our paper (synergylabs_spidergen).
@@
o1-preview $0.051\pm 0.029$ US $4.7\pm 0.28$
o1-mini $0.046\pm 0.023$ US $0.36\pm 0.03$
gpt-4o $0.039\pm 0.022$ US $0.64\pm 0.14$
Table 2: Average PMI Score and cost for SpiderGen using different LLMs across 40 product categories. SpiderGen performs better with models, such as OpenAI’s o1-preview and o1-mini. o1-mini provides a good tradeoff with good performance and lowest cost.
@@
We implemented SpiderGen using 3 OpenAI models: gpt-4o, o1-mini, and o1-preview. o1-mini and o1-preview are reasoning models, with o1-mini being a smaller, faster, and cheaper model than the other two. Table 2 compares the performance of these models in terms of the average normalized PMI and cost for generating a single PFG across 40 product categories. The two reasoning models (o1-mini and o1-preview) outperform gpt-4o since they are able to provide more detailed outputs. However, o1-mini is close in terms of performance to o1-preview while being 1/13th the cost at less than US \$1 on average per PFG. This cost is far lower than the typical production of a PFG, which can be upwards of US \$25000 (tasaki_international_2017).
@@
We evaluate the effect of producing a different number of sample products in the first step of SpiderGen (shown in Figure 3 for four product categories). We evaluate the production of PFGs using 3, 6, 9, 12, and 15 sample products for each PFG generation. As shown in the figure, the PMI often increases with the addition of more sample products. We note that less “niche” product categories, such as “Grain Mills” benefit from a larger number of sample products, and more “niche” categories, such as “Moka Coffee” benefit from smaller numbers of sample products, and can even have deteriorated PFGs, as the number of sample products increases. Since generating 15 products works well for many scenarios, we use it for our implementation of SpiderGen.
@@
To evaluate SpiderGen end-to-end, we execute our pipeline for 65 product categories and compare the generated PFGs to the ground truth graphs extracted from real LCA documents [^4] Due to limited available PCR documents, we solely utilize these 65 PCRs from EPD International (EPD) (EPD).
@@
We compared SpiderGen with the two baseline methods, LLMDirect and LLMExample, across 65 product categories. Our results show that SpiderGen exceeds both baseline methods, providing a median normalized PMI score of $0.043\pm 0.026$ across different product categories as compared to $0.026\pm 0.020$ for LLMDirect and $0.029\pm 0.023$ PMI for LLMExample. Based on our qualitative analysis of a subset of ten products, we find that SpiderGen avoids generating overly specific processes and captures more non-obvious auxiliary processes as compared to the baselines.
Limitations of LLMDirect: Although a Chain-of-Thought process may work for many problems, we found that an algorithm such as LLMDirect is insufficient to solve our problem. Firstly, it is unable to capture all the nuanced steps that are necessary for complete PFGs. LLMDirect results in frequently missed steps that are peripheral to the main manufacturing processes, such as downstream product maintenance and disposal steps. Further, PFG is sometimes overly specific to one subcategory of the product category.
Limitations of LLMExample: While LLMExample is more effective than LLMDirect, since it is able to capture the nuanced steps of the PFG by following an example, it is still insufficient to solve our problem. LLMExample misses peripheral steps, like “machine maintenance,” that may be important to a broader product category. Additionally, the PFG is still overly specific to a subset of items of the product category.
@@ 7 Discussion and Conclusion
In this work, we introduce SpiderGen, a novel machine learning workflow that automates the generation of PFGs, and introduce evaluation methods for measuring the correctness of PFGs. We present an implementation of SpiderGen, and provide an evaluation of SpiderGen across a set of 65 product categories. We find that SpiderGen is able to capture a large portion of required processes for tracking the carbon footprint of various product categories.
@@
Increasing transparency in LCA automation. Recent work indicates that providing transparency in automated LCA methods is a key criteria (ulissi_criteria_2025). Thus, in SpiderGen, we enable transparency by providing sample products, evidence and rationales for processes, and deterministic clustering techniques for forming coarser processes. We hope that this enables human LCA experts to gain a better understanding of how SpiderGen generates PFGs. However, more work must be done to achieve other evaluation goals, such as measuring and indicating the uncertainty of the PFG generation, as well as enabling LCA experts to collaboratively improve the PFG estimates. We hope to explore these possibilities in future work.
Enhancing System Boundaries of PFG Generation. We note that a primary limitation of SpiderGen is derived from the scope, or “system boundary”, of the PFG being in conflict with the system boundary defined appropriately for the product category. In the real world, the system boundary is human-defined. However, for our experiments, we made assumptions about the system boundaries based on ISO standards and encoded them into SpiderGen. We note that this is the root cause for SpiderGen missing processes or hallucinating processes. Although the precision of SpiderGen is relatively high (on average 81%), SpiderGen occasionally produces erroneous processes due to sticking closely to system boundaries that may be inappropriate for a given product category. For example, a more common error was to include processes such as “replacement”, or “quality control”, even if they were not relevant for a given product category. Additionally, SpiderGen misses some the auxiliary processes which are not directly related to the product manufacturing, material composition or usage.We believe that exploring human-AI collaboration methods to make appropriate choices for system boundaries is a promising next step for future work.
@@
Citing the Sources: Utilizing Retrieval Augmented Generation. In future iterations of SpiderGen, we believe that Retrieval Augmented Generation (RAG), where LLM models can search resources such as the internet, will be a promising method for increasing the transparency and traceability of PFG generation.
@@ Acknowledgments
This work is supported by the National Science Foundation Award CNS-2325956, as well as the National Science Foundation Graduate Research Fellowship Program under Grant DGE-2140739. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.